OpenStudy (anonymous):

The surface area of a cube is 100 square centimeters. Estimate to the nearest tenth of a centimeter the side length of the cube using the formula S = 6s2, where S represents surface area and s represents side length. Make sure to show all work and label your answer.

OpenStudy (compassionate):

Here at OpenStudy we lead the user to the answer. What do you think your first step should be?

OpenStudy (anonymous):

Lemme see....

OpenStudy (anonymous):

Make some kind of equation with what we know at the moment and that would be the formula

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Hello?

OpenStudy (compassionate):

Yes, yes! Find the length of one side first. 1,000/6 = 6/6s^2 166.66 166.7 = s^2 83.3 = s

OpenStudy (compassionate):

I have to go. Bye:)

OpenStudy (dan815):

compass is that ur kid?

OpenStudy (anonymous):

TY!

OpenStudy (dan815):

hey thats not the answer

OpenStudy (dan815):

lemme fix it for u, since he left already

OpenStudy (dan815):